Explain how Kirchhoff’s rules govern the voltage across and the current through each of the two resistors that are connected in series to a power supply. 


1
Expert's answer
2021-05-19T11:06:05-0400

According to Kirchhoff’s voltage law, the sum of sources of voltage is equal to the sum of potential drops across each resistor. Since the connection is series, the current is the same through both resistor:


"E=V_1+V_2=IR_1+IR_2=I(R_1+R_2)."